A leading magazine reported at one time that the average number of weeks an individual is unemployed is 39 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 39 weeks and that the population standard deviation is 3.3 weeks. Suppose you would like to select a random sample of 52 unemployed individuals for a follow-up study.

Find the probability that a single randomly selected value is greater than 39.6.
P(X > 39.6) = ______ (Enter your answers as numbers accurate to 4 decimal places.)

Find the probability that a sample of size n=52 is randomly selected with a mean greater than 39.6.
P(M > 39.6) = ______ (Enter your answers as numbers accurate to 4 decimal places.)
Answer #1

Solution : given that, u = 39 weeks. 6 = 3.3 weeks. p(x> 39.6) = 1-2/x. 39.6 - 39 3.3 = 1-P (2 2 0.18 ) = -0.5714 Probability